# NeMo-Aligner: Scalable Toolkit for Efficient Model Alignment

### v1

@@ 3.5 What contributes to system performance?
To better understand the importance of each aspect of NeMo-Aligner’s PPO system design, we conduct ablation studies by removing one aspect at a time and measuring the overall time per step as shown in Table 3. We find that TensorRT-LLM Integration is the most critical component for high system performance, without which PPO will take nearly seven times as long for each step. This is followed by using TensorRT Refit to avoid TensorRT-LLM engine recompiling (1.46x), the use of async requests between actor and critic models (1.29x) and finally load-balancing of data parallel workers during generation using a worker pool (1.08x).
@@
Optimal RLHF Setup 187.6 (13.8) 1x
- TensorRT-LLM Integration 1305.9 (160.9) 6.96x
@@
- TensorRT Refit 278.6 (7.8) 1.46x
- Async Requests 241.4 (16.6) 1.29x
- Worker Pool 203.1 (20.6) 1.08x
Table 3: Ablation studies on training Llama 2 70B actor and Llama 2 13B critic on Rollout batch size of 512 with 8 Actor nodes and 2 critic nodes. Nodes are 8*H100-80GB-SXM connected with intra-node NVLink (NVIDIA, 2023a) and inter-node Infiniband (NVIDIA, 2024) interconnects. Time per step is calculated based on mean of 5 steps after the first step, as the first step incurs substantial additional time for TRT-LLM Engine Building.
@@ Acknowledgements
Appendix A Appendix
No. of compute nodes (Actor + Critic) Tensor Parallel Pipeline Parallel Data Parallel
8 + 2 8 + 4 4 + 1 2 + 4
16 + 4 8 + 4 4 + 1 4 + 8
32 + 8 8 + 4 4 + 1 8 + 16
32 + 16 8 + 8 2 + 2 16 + 8
64 + 32 8 + 8 2 + 2 32 + 16
Table 5: Parallelism settings for scaling experiments.

### v2

@@ 3.5 What contributes to system performance?
Optimal RLHF Setup 53.7 (2.223) 1x
- TensorRT-LLM Integration 372.2 (37.433) 6.93x
@@
- Reshard 207.6 (8.940) 3.87x
- TensorRT Refit 169.0 (5.206) 3.15x
- Async Requests 82.8 (7.826) 1.54x
Table 4: Ablation studies on training Llama 3 70B actor and critic on rollout global batch size of 128 with 8 Actor nodes and 4 critic nodes. Nodes are 8*H100-80GB-SXM connected with intra-node NVLink (NVIDIA, 2023a) and inter-node Infiniband (NVIDIA, 2024) interconnects. Time per step is calculated based on mean of 5 steps after the first step, as the first step incurs substantial additional time for TRT-LLM Engine Building.
To better understand the importance of each aspect of NeMo-Aligner’s PPO system design, we conduct ablation studies by removing one aspect at a time and measuring the overall time per step as shown in Table 4. We find that TensorRT-LLM Integration is the most critical component for high system performance, without which PPO will take nearly seven times as long for each step. This is followed by resharding our model to use tensor parallelism only during inference (3.87x), using TensorRT Refit to avoid TensorRT-LLM engine recompiling (3.15x), the use of async requests between actor and critic models (1.54x). We did not observe meaningful speedup in using the worker pool to balance the work between data parallelism ranks because the problem size is small (with rollout global batch size of 128) and therefore the worker imbalance is less than the overhead we incur doing the balancing itself. Nevertheless, we expect this feature to help for larger problem sizes.
3.6 Training the largest open source LLM
No. of compute nodes (Actor + Critic) 84 + 42
Time per step in seconds (std.) $\downarrow$
Overall 164.60 (3.434)
   Train 5.60 (0.532)
   Rollout 158.90 (2.969)
    - Response generation 140.10 (1.067)
    - Log-probs calculation 17.20 (2.352)
    - TensorRT Refit 0.70 (0.108)
    - Critic wait 0.30 (0.454)
Table 5: NeMo Aligner supports the largest open source LLM as of July 2024 (Llama 3.1 405B) as both the actor and critic. We run the model on rollout global batch size of 128 and BF16 precision. Nodes are 8*H100-80GB-SXM connected with intra-node NVLink (NVIDIA, 2023a) and inter-node Infiniband (NVIDIA, 2024) interconnects. Time per step is calculated based on mean of 5 steps after the first step, as the first step incurs substantial additional time for TRT-LLM Engine Building. Further configuration details are in Table 3.
NeMo-Aligner supports alignment of the largest open source LLMs as of July 2024, such as Nemotron 4 340B (Nvidia et al., 2024) and Llama 3.1 405B (Dubey et al., 2024). In Table 5, we perform PPO on Llama 3.1 405B using 1008 H100 GPUs, using a configuration based on Sec. 3.3. We use Llama 3.1 405B Instruct as the actor and we train a Reward Model on top of Llama 3.1 405B Instruct using the same data and hyper-parameters as Nemotron 4 340B Reward (Wang et al., 2024). Compared to the 70B model, the 405B model is substantially slower to train during PPO, mainly bottle-necked by the slow response generation stage. This is because the 405B model generation cannot fit into a single node and requires pipeline-parallel generation which we plan to optimize further in future work. Additionally, in the first few steps of PPO, the 405B model generates longer responses (mean of 916 tokens) than the 70B model with a mean of 351 tokens.